# Linear complementary dual quasi-cyclic codes of index 2

### v2

@@ 5 Symplectic LCD quasi-cyclic codes of index 2
Assume that all the notations are the same as in the previous sections. First, we recall some definitions and results, for details see [11, Section V]. For $x=(x_{1}|x_{2})$ and $y=(y_{1}|y_{2})$ in $F^{2m}$, where $x_{i},y_{i}\in F^{m}$ for $i=1,2$, we have
@@
Let $C$ be a QC code of length $2m$ and index $2$ with CRT decomposition given in (2). By extending the map $\tau$ canonically to the vector spaces $(F_{i})^{2},(H_{j}^{\prime})^{2},(H_{j}^{\prime\prime})^{2}$ and applying the maps component-wise to (3), we obtain symplectic dual of $C$ (see [11, Proposition V.1]).
@@
Using the above characterization, we have the following characterization of symplectic LCD QC codes in terms of constituents (see [11, Eq. V.7]).
@@
Example 3.
Let $m=17$, $q=2$. Then $x^{m}-1=(x+1)(x^{8}+x^{5}+x^{4}+x^{3}+1)(x^{8}+x^{7}+x^{6}+x^{4}+x^{2}+x+1)\in\mathbb{F}_{2}[x]$. Take
$\begin{split}g_{11}=&x+1,\\ g_{22}=&(x+1)(x^{8}+x^{7}+x^{6}+x^{4}+x^{2}+x+1),\\ g_{12}=&(x+1)(x^{4}+x^{3}+x^{2}+x+1).\end{split}$
Then the code generated by $(g_{11},g_{12})$ and $(0,g_{22})$ is a symplectic LCD code with parameters $[34,24,4]$. It is optimal as per Grassl’s table.
@@
This result is consistent with that in [14].
Restricting to the case $r_{11}(x)=r_{22}(x)=1$ in Theorem 5.3, we can get the following sufficient condition for a code to be symplectic LCD.
@@
In the following example, we construct a quasi-cyclic (symplectic) LCD code using Corollary 5.7.
Example 4.
Let $m=15$, $q=2$. Then $x^{m}-1=(x+1)(x^{2}+x+1)(x^{4}+x+1)(x^{4}+x^{3}+1)(x^{4}+x^{3}+x^{2}+x+1)\in\mathbb{F}_{2}[x]$. Take
$\begin{split}g_{11}=&(x^{4}+x^{3}+x^{2}+x+1)(x^{4}+x+1),\\ g_{22}=&(x^{4}+x^{3}+x^{2}+x+1)(x^{4}+x^{3}+1),\\ g_{12}=&x^{4}+x^{3}+x^{2}+x+1.\end{split}$
Then the code generated by $(g_{11},g_{12})$ and $(0,g_{22})$ is a symplectic LCD code with parameters $[30,14,5]$.
@@ 6 Hermitian LCD quasi-cyclic codes of index 2
The constituent code $C_{j}^{\prime}$ over $H_{j}^{\prime}$ is obtained by evaluating $a(x)$ at $\xi^{v_{j}}$, while the constituent code $C_{j}^{\prime\prime}$ over $H_{j}^{\prime\prime}$ is obtained by evaluating $\hat{a}(x)$ at $\xi^{v_{j}}$. In [11], the authors utilized a slightly different isomorphism. They defined the constituent code $C_{j}^{\prime}$ over $H_{j}^{\prime}$ by evaluating $a(x)$ at $\xi^{v_{j}}$ and the constituent code $C_{j}^{\prime\prime}$ over $H_{j}^{\prime\prime}$ by evaluating $a(x)$ at $\xi^{-qv_{j}}$, which means $a(\xi^{-qv_{j}})=(\hat{a}(\xi^{v_{j}}))^{q}$ (see [11, Eq. (IV.10)] and remark after Eq. (IV.14)).
@@
We have the following characterization of Hermitian LCD QC codes in terms of constituents (see [11, Eq. (IV.16)]).
@@
Next, using Theorem 6.3, we construct several new and optimal quasi-cyclic Hermitian LCD codes over $\mathbb{F}_{4}=\mathbb{F}_{2}[w]$ with $w^{2}+w+1=0$ (in Table 2).
@@
$[22,11,8]$ $x+1$ $(x+1)$ $(w^{2}x+wx^{2}$ $x^{4}+wx^{6})$ $(x^{5}+wx^{4}+x^{3}$ $x^{2}+w^{2}x+1)$ $(x^{5}+w^{2}x^{4}+x^{3}$ $x^{2}+wx+1)$ BKLC
$[22,16,4]$ $x+1$ $x^{5}+wx^{4}$ $x^{2}+x+1$ $x^{5}+wx^{4}+x^{3}$ $x^{2}+w^{2}x+1$ Optimal Hermitian LCD
$[26,13,8]$ $x+1$ $w^{2}x^{9}+x^{8}+w^{2}x^{7}$ $x^{6}+x^{4}+wx^{3} +$$x^{2}+wx$ $(x^{6}+wx^{5}+w^{2}x^{3}+wx+1)$ $(x^{6}+w^{2}x^{5}+wx^{3}+w^{2}x+1)$ -
$[34,16,11]$ $(x+1)(x^{4}+wx^{3}$ $x^{2}+wx+1)$ $(x+1)$ $(x^{9}+w^{2}x^{8}+x^{7}$ $wx^{2}+w^{2}x)$ $(x+1)$ $(x^{4}+x^{3}+wx^{2}+x+1)$ $(x^{4}+x^{3}+w^{2}x^{2}+x+1)$ $(x^{4}+w^{2}x^{3}+x^{2}+w^{2}x+1)$ New
$[34,17,9]$ $(x+1)(x^{4}+w^{2}x^{3}$ $x^{2}+w^{2}x+1)$ $x^{10}+x^{8}+w^{2}x^{7}$ $w^{2}x^{6}+x^{2}+x$ $(x^{4}+x^{3}+wx^{2}+x+1)$ $(x^{4}+x^{3}+w^{2}x^{2}+x+1)$ $(x^{4}+wx^{3}+x^{2}+wx+1)$ -
Table 2: Quaternary quasi-cyclic (Hermitian) LCD codes, where BKLC means that the minimum distance of the code is equal to the minimum distance of Best Known Linear Code according to the Grassl’s table, “New” means codes are new compare to papers [14],[20],[42] and “Optimal Hermitian LCD” means there exists no LCD code with parameters $[22,16,5]$ (see [31]).
@@ Conclusion
In this work, we have given a nice polynomial-based characterization of quasi-cyclic linear complementary dual (LCD) codes of index $2$ with respect to the Euclidean, Hermitian and symplectic inner products. Using these characterizations, we constructed several binary, ternary, and quaternary LCD codes. Our results extend the existing characterizations obtained for one-generator quasi-cyclic codes. Moreover, the techniques introduced in our characterization can be readily generalized to the broader class of quasi-twisted codes. A promising direction for future research is to obtain polynomial characterization for quasi-cyclic LCD codes of arbitrary index. This extension would require handling a large number of polynomials together, which will be more complicated.

### v3

@@ 5 Symplectic LCD quasi-cyclic codes of index 2
Assume that all the notations are the same as in the previous sections. First, we recall some definitions and results, for details see [15, Section V]. For $x=(x_{1}|x_{2})$ and $y=(y_{1}|y_{2})$ in $F^{2m}$, where $x_{i},y_{i}\in F^{m}$ for $i=1,2$, we have
@@
Let $C$ be a QC code of length $2m$ and index $2$ with CRT decomposition given in (2). By extending the map $\tau$ canonically to the vector spaces $(F_{i})^{2},(H_{j}^{\prime})^{2},(H_{j}^{\prime\prime})^{2}$ and applying the maps component-wise to (3), we obtain symplectic dual of $C$ (see [15, Proposition V.1]).
@@
Using the above characterization, we have the following characterization of symplectic LCD QC codes in terms of constituents (see [15, Eq. V.7]).
@@
This result is consistent with that in [19]. Restricting to the case $r_{11}(x)=r_{22}(x)=1$ in Theorem 5.3, we can get the following sufficient condition for a code to be symplectic LCD.
@@
We now describe some examples of binary quasi-cyclic symplectic LCD codes using Theorems 5.3, 5.6 and Corollary 5.7. In [8], Calderbank et al. proved that the binary symplectic inner product and the quaternary trace-Hermitian inner product are equivalent. Therefore, a binary symplectic LCD code with parameters $2n,k,d_{s}]_{2}$ is also a quaternary additive complementary dual (ACD) code with parameters $(n,k/2,d_{s})_{4}$.
We note that it is still a challenging problem to efficiently construct quaternary additive codes that outperform linear codes, see [20]. In Table 5, we describe quatenary ACD codes which match the performance of Hermitian LCD codes in [4] and ACD codes in [20].
ACD code $g_{11}(x)$ $g_{12}(x)$ $g_{22}(x)$ Comparison
$(15,5,8)_{4}$ $(x+1)(x^{4}+x^{3}+x^{2}$ +x+1)+\,x+1) $x(x+1)(x^{4}+x^{3}+x^{2}+x+1)$ $(x^{3}+x+1)$ $x^{15}-1$ $15,5,8]_{4}^{H}$, see [4]
$(15,6,7)_{4}$ $x^{4}+x^{3}+x^{2}+x+1$ $(x+1)(x^{3}+x^{2}+1)$ $(x^{4}+x^{3}+x^{2}+x+1)$ $(x^{4}+x^{3}+x^{2}+x+1)(x^{4}+x^{3}+1)$ $(x^{4}+x+1)(x^{2}+x+1)$ $15,6,7]_{4}^{H}$, see [4]
$(15,9,5)_{4}$ 11 $(x+1)^{2}(x^{4}+x+1)$ $(x^{4}+x+1)(x^{4}+x^{3}+1)$ $(x^{4}+x^{3}+x^{2}+x+1)$ $15,9,5]_{4}^{H}$, see [4]
$(19,9,8)_{4}$ $x+1$ $(x+1)^{3}(x^{4}+x^{3}+x^{2}+x+1)$ $(x^{11}+x^{6}+x^{5}+x^{4}+1)$ $x^{19}-1$ $19,9,8]_{4}^{H}$, see [4]
$(19,10,7)_{4}$ 11 $x^{11}+x^{8}+x^{7}+x^{6}$ $x^{4}+x^{3}+x^{2}+x$ $x^{18}+x^{17}+x^{16}+x^{15}+x^{14}+x^{13}$ $x^{12}+x^{11}+x^{10}+x^{9}+x^{8}+x^{7} +$$x^{6}+x^{5}+x^{4}+x^{3}+x^{2}+x+1$ $19,10,7]_{4}^{H}$, see [4]
$(25,13,8)_{4}$ 11 $x^{23}+x^{12}+x^{11}+x^{10}$ $x^{9}+x^{5}+x$ $(x^{4}+x^{3}+x^{2}+x+1)$ $(x^{20}+x^{15}+x^{10}+x^{5}+1)$ $25,13,8]_{4}^{H}$, see [4], [20]
$(27,13,9)_{4}$ $x^{2}+x+1$ $x^{2}(x+1)^{3}(x^{2}+x+1)^{3}$ $(x^{7}+x+1)$ $(x^{2}+x+1)$ $(x^{6}+x^{3}+1)(x^{18}+x^{9}+1)$ $27,13,9]_{4}^{H}$, see [4]
Table 5: Additive complementary dual codes from Theorems 5.3, 5.6 and Corollary 5.7.
@@ 6 Hermitian LCD quasi-cyclic codes of index 2
The constituent code $C_{j}^{\prime}$ over $H_{j}^{\prime}$ is obtained by evaluating $a(x)$ at $\xi^{v_{j}}$, while the constituent code $C_{j}^{\prime\prime}$ over $H_{j}^{\prime\prime}$ is obtained by evaluating $\hat{a}(x)$ at $\xi^{v_{j}}$. In [15], the authors utilized a slightly different isomorphism. They defined the constituent code $C_{j}^{\prime}$ over $H_{j}^{\prime}$ by evaluating $a(x)$ at $\xi^{v_{j}}$ and the constituent code $C_{j}^{\prime\prime}$ over $H_{j}^{\prime\prime}$ by evaluating $a(x)$ at $\xi^{-qv_{j}}$, which means $a(\xi^{-qv_{j}})=(\hat{a}(\xi^{v_{j}}))^{q}$ (see [15, Eq. (IV.10)] and remark after Eq. (IV.14)).
@@
We have the following characterization of Hermitian LCD QC codes in terms of constituents (see [15, Eq. (IV.16)]).
@@
In [11], Carlet et al. showed that any linear code is equivalent to a Hermitian LCD code for $q>4$. This motivates us to search for quaternary Hermitian LCD codes with good parameters. In Tables 6 and 7, we provide several examples of quasi-cyclic Hermitian LCD codes over $\mathbb{F}_{4}=\mathbb{F}_{2}[w]$ with $w^{2}+w+1=0$, using Theorem 6.3.
In Table 6, we use the abbreviation BKLC to indicate that a code is the best known linear code according to [17]. Similarly, BKHLCD denotes the best known Hermitian LCD code, according to [4] and [46].
In Table 7, we present examples of Hermitian LCD codes with minimum distance larger than previously known codes in the literature [13, 18] with the same length and dimension. The codes reported in [13, 18] were used as benchmarks for ACD codes in [52].
@@
$4,8,5]_{4}^{H} 11$ $wx^{5}+w^{2}x^{4}+wx^{3}$ $x^{2}+x+1$ $(x^{3}+x+1)$ $(x^{3}+x^{2}+1)$ $d=d_{\text{BKHLCD}}$, see [4]
$22,11,8]_{4}^{H}$ $x+1$ $(x+1)$ $(w^{2}x+wx^{2}+\,x^{4}+wx^{6})$ $(x^{5}+wx^{4}+x^{3}+x^{2}+w^{2}x+1)$ $(x^{5}+w^{2}x^{4}+x^{3}+x^{2}+wx+1)$ $d=d_{\text{BKHLCD}}$, see [4]
$22,12,7]_{4}^{H}$ 11 $w^{2}x^{7}+w^{2}x^{5}+w^{2}x^{4}$ $w^{2}x^{2}+w^{2}x+1$ $(x^{5}+wx^{4}+x^{3}+x^{2}+w^{2}x+1)$ $(x^{5}+w^{2}x^{4}+x^{3}+x^{2}+wx+1)$ $d=d_{\text{BKHLCD}}$, see [4]
$22,16,4]_{4}^{H}$ $x+1$ $x^{5}+wx^{4}$ $x^{2}+x+1$ $x^{5}+wx^{4}+x^{3}+x^{2}+w^{2}x+1$ $d=d_{\text{BKHLCD}}$, see [4]
$22,17,4]_{4}^{H}$ 11 $w^{2}x^{4}+x^{2}+x+w$ $x^{5}+wx^{4}+x^{3}+x^{2}+w^{2}x+1$ $d=d_{\text{BKHLCD}}$, see [4]
$26,13,8]_{4}^{H}$ $x+1$ $w^{2}x^{9}+x^{8}+w^{2}x^{7}$ $x^{6}+x^{4}+wx^{3} +$$x^{2}+wx$ $(x^{6}+wx^{5}+w^{2}x^{3}+wx+1)$ $(x^{6}+w^{2}x^{5}+wx^{3}+w^{2}x+1)$ $d=d_{\text{BKHLCD}}$, see [4]
$34,16,11]_{4}^{H}$ $(x+1)(x^{4}+wx^{3}$ $x^{2}+wx+1)$ $(x+1)$ $(x^{9}+w^{2}x^{8}+x^{7}$ $wx^{2}+w^{2}x)$ $(x+1)$ $(x^{4}+x^{3}+wx^{2}+x+1)$ $(x^{4}+x^{3}+w^{2}x^{2}+x+1)$ $(x^{4}+w^{2}x^{3}+x^{2}+w^{2}x+1)$ $d=d_{\text{BKLC}}-1$, see [17]
$34,17,9]_{4}^{H}$ $(x+1)(x^{4}+w^{2}x^{3}$ $x^{2}+w^{2}x+1)$ $x^{10}+x^{8}+w^{2}x^{7}$ $w^{2}x^{6}+x^{2}+x$ $(x^{4}+x^{3}+wx^{2}+x+1)$ $(x^{4}+x^{3}+w^{2}x^{2}+x+1)$ $(x^{4}+wx^{3}+x^{2}+wx+1)$ $d=d_{\text{BKHLCD}}$, see [46]
$34,24,6]_{4}^{H}$ $x+1$ $(x+1)(x+w^{2})(x^{2}+wx+w)$ $(x^{4}+x^{3}+x+w^{2})$ $(x+1)$ $(x^{4}+wx^{3}+x^{2}+wx+1)$ $(x^{4}+w^{2}x^{3}+x^{2}+w^{2}x+1)$ $d=d_{\text{BKLC}}$, see [17]
$34,26,5]_{4}^{H}$ 11 $x^{7}+wx^{5}+w^{2}x^{4}+x^{3}$ $w^{2}x^{2}+w$ $(x^{4}+wx^{3}+x^{2}+wx+1)$ $(x^{4}+w^{2}x^{3}+x^{2}+w^{2}x+1)$ $d=d_{\text{BKLC}}$, see [17]
$38,29,5]_{4}^{H}$ 11 $w^{2}x^{8}+w^{2}x^{7}+w^{2}x^{6}$ $wx^{5}+wx+1$ $x^{9}+wx^{8}+wx^{6}+wx^{5}$ $w^{2}x^{4}+w^{2}x^{3}+w^{2}x+1$ $d=d_{\text{BKLC}}-1$, see [17]
Table 6: Quaternary quasi-cyclic Hermitian LCD codes from Theorem 6.3, length $14~\leq~n~\leq~38$.
$C$ $g_{11}(x)$ $g_{12}(x)$ $g_{22}(x)$ Parameters from [13, 18]
$46,23,8]_{4}^{H}$ $x+1$ $w^{2}x^{21}+x^{20}+wx^{15}$ $x^{14}+x^{8}+w^{2}x^{2}$ $(x^{11}+x^{9}+x^{7}+x^{6}+x^{5}+x+1)$ $(x^{11}+x^{10}+x^{6}+x^{5}+x^{4}+x^{2}+1)$ $46,23,7]_{4}^{H}$
$58,29,8]_{4}^{H}$ $x+1$ $x^{27}+w^{2}x^{19}+w^{2}x^{15}$ $x^{14}+x^{12}+wx^{10}$ $x^{28}+x^{27}+x^{26}+x^{25}+x^{24}+x^{23}+x^{22}+x^{21}$ $x^{20}+x^{19}+x^{18}+x^{17}+x^{16}+x^{15}+x^{14}+x^{13} +$$x^{12}+x^{11}+x^{10}+x^{9}+x^{8}+x^{7}+x^{6}+x^{5} +$$x^{4}+x^{3}+x^{2}+x+1$ $58,29,4]_{4}^{H}$
$62,31,8]_{4}^{H}$ $x+1$ $x^{22}+x^{18}+x^{14}+wx^{12}$ +x+1+\,x+1 $(x^{5}+x^{2}+1)(x^{5}+x^{3}+1)(x^{5}+x^{3}+x^{2}+x+1)$ $(x^{5}+x^{4}+x^{2}+x+1)(x^{5}+x^{4}+x^{3}+x+1)$ $(x^{5}+x^{4}+x^{3}+x^{2}+1)$ $62,31,7]_{4}^{H}$
$70,35,8]_{4}^{H}$ $x+1$ $w^{2}x^{25}+wx^{22}+x^{19}+x^{18}$ $w^{2}x^{17}+x^{16}$ $(x^{3}+x+1)(x^{3}+x^{2}+1)(x^{4}+x^{3}+x^{2}+x+1)$ $(x^{12}+x^{10}+x^{9}+x^{8}+x^{7}+x^{4}+x^{2}+x+1)$ $(x^{12}+x^{11}+x^{10}+x^{8}+x^{5}+x^{4}+x^{3}+x^{2}+1)$ $70,35,5]_{4}^{H}$
$74,37,8]_{4}^{H}$ $x+1$ $x^{29}+w^{2}x^{27}+w^{2}x^{24}$ $x^{8}+wx^{5}+1$ $x^{36}+x^{35}+x^{34}+x^{33}+x^{32}+x^{31}+x^{30}+x^{29}$ $x^{28}+x^{27}+x^{26}+x^{25}+x^{24}+x^{23}+x^{22}+x^{21} +$$x^{20}+x^{19}+x^{18}+x^{17}+x^{16}+x^{15}+x^{14}+x^{13} +$$x^{12}+x^{11}+x^{10}+x^{9}+x^{8}+x^{7}+x^{6}+x^{5} +$$x^{4}+x^{3}+x^{2}+x+1$ $74,37,4]_{4}^{H}$
Table 7: Quaternary quasi-cyclic Hermitian LCD codes from Theorem 6.3, length $46~\leq~n~\leq~74$.
@@ Conclusion
In this work, we have provided polynomial-based characterizations of quasi-cyclic linear complementary dual (LCD) codes of index $2$ with respect to the Euclidean, Hermitian and symplectic inner products. Using these characterizations, we constructed several binary, ternary, and quaternary LCD codes. Our results extend the existing characterizations obtained for one-generator quasi-cyclic codes. Moreover, the techniques introduced in our characterization can be readily generalized to the broader class of quasi-twisted codes. A promising direction for future research is to obtain polynomial characterization for quasi-cyclic LCD codes of arbitrary index. This extension would require handling a large number of polynomials together, which will be more complicated.